OpenStudy (anonymous):

3+sqrt3/3-sqrt3(3+sqrt3/3+sqrt3) sorry it looks confusing, I'll fix it

OpenStudy (anonymous):

\[\frac{ 3+\sqrt{3} }{ 3-\sqrt{3} } (\frac{ 3+\sqrt{3} }{ 3+\sqrt{3} } )\]

OpenStudy (anonymous):

pretty sure the next step is 9+6sqrt3+3/9-3 but I don't know how to get that

OpenStudy (anonymous):

I think I'm supposed to foil it, but I dunno

OpenStudy (anonymous):

(9+3+6sqrt3)/(9-3) =2 +sqrt3

OpenStudy (anonymous):

so you're applying the conjugate... \[\frac{ (3+\sqrt{3})(3+\sqrt{3)} }{ 9-3 }\]

OpenStudy (anonymous):

(a+b)^2 = a2 + b2 +2ab (a+b)(a-b)= a2 - b2

OpenStudy (anonymous):

you are right @koel I'm just having trouble doing the algebra part

OpenStudy (anonymous):

3x3 is 9, I get that, but where do the 6 and 3 come from?

OpenStudy (anonymous):

(3+sqrt3)^2 = 9 +6 sqrt3 +3

OpenStudy (anonymous):

and (3 -sqrt 3)(3+sqrt 3) = 9-3 =6

OpenStudy (anonymous):

ok, I kinda see... 3^2 is 9... does the sqrt3^2 = 6?

OpenStudy (anonymous):

no sqrt3^2 = 3

OpenStudy (anonymous):

ok, where does the 6 come from then?

OpenStudy (anonymous):

its the 2ab term because of which 6sqrt3 comes

OpenStudy (anonymous):

I'm trying really hard to remember how that equation works. it's been a while. what is it called?

OpenStudy (anonymous):

the (a+b)^2...

OpenStudy (anonymous):

(a+b)^2 = a^2 + 2ab + b^2 a =3 b=sqrt3 so a^2=9 2ab = 2*3*sqrt3 = 6sqrt3 b^2 = sqrt3^2 =3

OpenStudy (anonymous):

i dont think the formula has a name you can derive it by multiplying (a+b)(a+b) = a(a+b)+b(a+b)

OpenStudy (anonymous):

oh yea! ok, lemme try that